# Neutron star evolution by combining discontinuous Galerkin and finite volume methods

### v1

@@ IV.2 Special relativistic blast wave tests
As a next step, we propagated a so-called special relativistic blast wave in one dimension. The initial setup has of discontinuities in the matter fields of $\rho_{0}$ and $P$, as in a Riemann problem. We study the case from [120] with $(\rho_{0},P,v_{x})$ given by
@@
We set up the mesh with $16$ adjacent Cartesian elements along the $x$-axis. Initially, we set the element at the left boundary and the middle element containing the initial discontinuity to FV/FD containing $24$ equally spaced grid points, and the rest of the elements to DG, with $12$ grid points. During the evolution, we use our TEI from Sec. III.7 to decide if an element should use the DG or FV/FD method. For this case, we switch off the criterion from Eq. (50), as the purpose of it is to handle low density regions present in neutron star simulations, which is not relevant in this simulation. We also switch off the RDMP criterion, as it activates in regions where the analytic solution has zero slope, due to small amplitude variations in the numerical solution. These small amplitude variations can also trigger the Persson indicator if not relaxed enough. So, for the condition from Eq. (52), we choose $\alpha_{N}^{\mathrm{evo}}=4$ for the DG elements and $\alpha_{N}^{\mathrm{evo}}=5$ for the FV/FD elements. These values are lower than what we use for our star simulations in Secs. IV.3.2, IV.5.3 and IV.6, where we choose more conservative ones that result in more elements using FV/FD.
In Fig. (4) we plot the primitive variables. We show the entire mesh (top plot) as well as the contact discontinuity and the shock front (bottom plot), after evolving until time $t=0.4$. We see that the elements near the non-smooth regions evolve with FV/FD, while most other elements evolve with DG, as can be seen from the top plot. In both plots, we see that the fields in our simulation agree very closely with the analytic solutions that are plotted as solid black lines. This demonstrates that our hybrid DG+FV/FD method can successfully handle evolutions of shock fronts in matter fields, which is a crucial initial test, as such hydrodynamic shocks are known to develop during late inspiral and merger phases of a binary neutron star merger.
@@ IV.3.1 Pure DG (S) simulations:
For the DG case, we set up the mesh using a cubical patch at the center, surrounded by $6$ cubed sphere patches on all sides, which in turn have another layer of $6$ cubed sphere patches forming a spherical shell covering them, resulting in total 13 patches. We then h-refined the central cubical patch successively $4$, $5$ and $6$ times, while refining all the other patches one time less in all the cases, resulting in $10240$, $81920$, and $655360$ total elements after h-refining. In each of these elements, we use $5\times 5\times 5$ grid points. The star is contained in the central cube with side length $32$, which is centered on the star, while the outermost boundary of the entire mesh is at 320. The resultant initial number of elements across the star for the $l=4,5$, and $6$ are stated in Table 2. We set the time steps for the different resolutions to be $0.1$, $0.05$ and $0.025$ respectively. For the artificial atmosphere, we set $f_{\mathrm{atmo,level}}=10^{-12}$, $f_{\mathrm{atmo,cut}}=1.01\times 10^{-12}$, $f_{\mathrm{atmo,max}}=10$, $\epsilon_{\mathrm{atmo,max}}=0.1$, and $(Wv)_{\mathrm{atmo,max}}=0.01$ as per the description in Sec. III.5.1. We also apply our positivity limiters with elementwise limiting of $D$, $\tau$ and $S_{i}$, but not the pointwise fixes described in Sec. III.5.2.
@@
Figure 6: This figure shows the Fourier spectrum of the central point density $D_{\mathrm{C}}$ for (S) and (P) with pure DG for the simulation with $l=5$ levels of h-refinement for each case. The dashed vertical lines correspond to the expected frequencies for the various oscillations modes of the star, as stated in [121]. We see that the frequency for the principal mode from the oscillations matches perfectly with the theoretical value [121]. The higher modes are not obtained as pure DG simulations have too much noise due to Gibbs oscillations.
@@
This limiter treatment is essentially the same as that described in [76]. We restrict the lower limit of $D$ to $10^{-12}\rho_{0,\mathrm{max}}(t=0)$ and $\tau$ to 00, while enforcing $S<D+\tau$, as described when limiting. For the matter fields, as per Eq. (44), the filter uses $s=32$, while we use $s=40$ for the GH system. For both cases we have $\alpha_{f}=36$. For these simulations, we do not use the conserved to primitive variable conversion prescription mentioned in Sec. III.6. We retain the method for conversion described in Sec. 4.3.1 of [76]. This is because we performed simulation for the same physical system successfully under Cowling approximation with this method that has been presented in Sec. 4.3.4 in [76]. For the GH system we set the constraint damping parameters to be $\gamma_{0}=0.1$, $\gamma_{1}=-1$ and $\gamma_{2}=1$. The gauge is set such that the gauge source function is initialized according to the initial metric and then is kept constant throughout the simulation, i.e, we set
@@
[…]
@@
The run with $5$ times h-refined central cube was evolved stably until a time of $t=5000M_{\odot}$, even though the plots for this case only show data until $t=2000M_{\odot}$ time due to high computational costs. As can be seen from Fig. 5, the central density oscillates with time, with higher erratic oscillations towards the beginning. These errors are larger in the beginning in this pure DG simulation due to strong Gibbs oscillations arising from the surface of the star. This is because in the initial setup, the non-smooth feature on the surface of the star is at its sharpest, which produces large Gibbs oscillation in the DG scheme, as the surface lies in the middle of elements. These oscillations spread throughout the star and produce strong numerical disturbances. However, as the simulation progresses, our stabilizing mechanisms such as the positivity limiters, filters and the atmosphere damp the spurious Gibbs oscillations, which leads to smoother star oscillations. Hence, at later times, we see a systematic oscillation for the two higher resolutions, which is in agreement with the spectrum expected from such a setup, as can be seen from Fig. 6, that shows the Fourier spectrum of the central density $D_{C}(t)$ for the case with $5$ central refinement levels for (S) with pure DG. In this figure we see that the principal frequency matches the predicted value from literature [121]. The higher modes are buried in numerical noise due to the lack of resolution. Note that the numerical noise in the oscillations decreases with increasing resolution. We see the highest resolution is demonstrating consistent and persistent oscillations at the proper frequency, while for the middle resolution, they get damped slowly by the filters we apply. The lowest resolution of $4$ central refinement levels is swamped by Gibbs oscillations, and is depicted only to show improvement with resolution. We next consider mass conservation by looking at the volume integral of the rest mass density $D$ shown in the left plot of Fig. 7. It is very clear that mass conservation improves drastically with increasing resolution. The small amount of discrepancy in the highest resolution is attributed to the fact that we use positivity limiters that conserve the mass only when the element average is above the limits imposed. Furthermore, pointwise corrections to $D$ arise due to the atmosphere treatment. In this case, the $D$ value is pushed up whenever it wants to fall below the floor value set by us. Hence, in these situations the integral can increase.
@@
We have also performed strong scaling tests with this (S) setup with pure DG, to demonstrate Nmesh’s capability to scale efficiently on a large number of computational cores or processors. This is shown in Fig. 8. This plot shows speed up vs. number of cores used for the run. Speed up when using $N$ number of compute cores is defined as $t_{r}(1)/t_{r}(N)$, where $t_{r}(N)$ is the run time measured for $N$ cores, with each core running one MPI process. Since a single MPI process cannot obtain enough memory for the simulations, $t_{r}(1)$ is estimated from $t_{r}(1)=N_{\mathrm{min}}t_{r}(N_{\mathrm{min}})$, where $N_{\mathrm{min}}$ is for the run with the lowest number of MPI processes performed in each case.
We performed these runs on both the Bridges-2 and Expanse supercomputers. In our case, we have $N_{\mathrm{min}}=3$ for the Expanse supercomputer and $512$ for the Bridges-2 supercomputer, which is used for the data in Fig. 8. We go up to $4096$ cores on Expanse and $6400$ cores on Bridges-2, which were the maximum core count available to us in the respective computers. The simulation we tested this with had total $262144$ elements in the mesh. We plot both the real simulation run time data and the ideal ones for the number of cores we have in each of the cases, and show the speedups for both. The ideal speed-ups are calculated from the $t_{r}(N_{\mathrm{min}})$, as is the usual custom in such scaling estimates. We can see that we obtain good strong scaling, as the real speedups not only rise almost in a straight line, but track the ideal lines very closely. The fact that they overshoot somewhat is because we have only an estimate from $t_{r}(1)$ from $t_{r}(N_{\mathrm{min}})$. This confirms that Nmesh scales efficiently with an increasing number of cores or MPI processes.
Figure 8: This plot shows speedup data for strong scaling test on Bridges-2 and Expanse for the pure DG (S) simulation with $262144$ elements. The dashed lines represent ideal speedup curves extended till the range of CPUs used for each supercomputer, while the solid lines represent the actual measured data for Nmesh. As we can see, the ideal and real data are in good agreement, demonstrating strong scaling.
IV.3.2 DG+FV/FD hybrid (S) simulation:
@@
We also evolve (S) with our hybrid DG+FV/FD scheme. In the elements that use FV/FD, we use the WENOZ2 reconstruction scheme as described in Sec. III.2. Before using combined DG+FV/FD method, we have performed tests with pure FV/FD with the WENOZ2 to validate our new method. However, we chose case (U) for these instead of (S), as it is significantly more challenging, and hence offers much more robust and general conclusions. These results are presented in Sec. IV.5.1 and IV.5.2. For the WENOZ2 method in this case, we turn on extrapolation near the boundary as per the $w=4/3$ setup for Eq. III.2 described in Sec. III.2. Hence, we also switch on the cutoff levels for using $w=1$ as when $\rho_{0}<10^{-9}\rho_{0,max}(t=0)$ or $\epsilon<10^{-9}\epsilon_{max}(t=0)$, as described there. When the surface points of neighboring elements do not coincide, we use parabolic interpolation to exchange data between neighboring elements at the boundary. In the FD scheme used for the GH system we use fourth order dissipation in the interior of the elements (see Eq. (33)) with $\sigma=0.1$. Since we want the high order WENOZ used in the interior of elements to dominate the result, we increase the number of points in each direction in each FV/FD element to $n=16$, while in the DG elements, we use half as many points ($n=8$), as DG can provide high accuracy with lower number of points in the smooth regions. Each element then has $n\times n\times n$ total points. For this simulation, we set up the mesh similar to what we use for case (B), which is described later in Sec. IV.6, as they are the same star, with and without an imparted boost. The detail of the initial mesh can be seen from Fig. 9, which shows plots on the $xy$-plane. In the left top plot of the figure, we see the zoomed in view of the star. Initially, we set FV/FD only at the troubled region at the edge of the star. In order to determine which elements should be set to FV/FD at the initial time, we look at the coefficients of $\rho_{0}$ in an expansion in Legendre polynomials, as described in Sec. 2.5 of [76]. As we go from lower to higher order coefficients, we find a much slower drop off for elements that contain the star surface, than for elements in the star interior. We use this drop off to initially set only the elements containing the star surface to use FV/FD. These are recognizable from the denser grid lines as FV/FD uses twice as many points as DG. The initial number of DG and FV/FD elements across the star is given in Table 2. Next, in the top right plot we see the central cubical region, which is the best resolved region of the mesh. Here the size of this cube is $62$. However, as we can see, the resolution is not uniform. This is because we do not apply uniform h-refinement to the entire mesh. We h-refine a spherical region of radius $15$ centered on the star to $l=5$ times, which is the highest refined region in the mesh. The h-refinement is then dropped by one level as we move away from the star, with the different h-refinement spheres having radii $15,30,60$ and $120$. These successive regions can be seen in the plots on the top right, bottom left and bottom right. In the left bottom picture, we see the mesh up to the first cubed sphere patches surrounding the central cube, up to a distance of $104$. The last bottom right picture shows the entire mesh, which has a radius of $208$. In total, we have 13 patches similar to the setup in Sec. IV.3.1. The time step is set as per Sec. III.3.
@@
For the atmosphere we use the values $f_{\mathrm{atmo,level}}=10^{-11}$, $f_{\mathrm{atmo,cut}}=10^{-9}$, and $f_{\mathrm{atmo,max}}=1$, as this is the setup we also use for (U) and (B) simulations with DG+FV/FD, which are described later. The other atmosphere limits are still maintained at $\epsilon_{\mathrm{atmo,max}}=0.1$, and $(Wv)_{\mathrm{atmo,max}}=0.01$. Here, we employ the entire limiting process described in Sec. III.5.2 for $D$, $\tau$ and $S_{i}$, including both elementwise and pointwise limiting. For this purpose, we set $f_{S}^{\mathrm{lim}}=0.99$. In this simulation, we use the conversion procedure from conserved to primitive variables stated in Sec. III.6, setting $\epsilon_{\mathrm{floor}}=0.0$. The constraint damping parameters for the GH system are set to be the position dependent functions of the form from Eq. (5) along with Eqs. (59)-(61). The gauge is chosen to be the harmonic gauge, where the gauge source function is set as $H_{\alpha}=0$. In the DG elements, we filter the conserved matter variables $D,\tau$ and $S_{i}$ with $s=40$, while using $s=64$ for the GH variables of the evolved metric and its derivatives. We maintain $\alpha_{f}=36$ for both. No filters are used in the FV/FD elements, as dissipation plays the counterpart role. When incorporating the TEI described in Sec. III.7, we set the value of $\alpha_{N}^{\mathrm{evo}}$ from Eq. (52) in the elements using DG to be $8$ and to $9$ in the elements using FV/FD. For this criterion, we only consider the coefficients that are greater than $f_{\mathrm{unfiltered}}=0.9$ times their unfiltered value. We set the cutoff from Eq. (50) as $\rho_{\mathrm{cut,low}}=10^{-4}\rho_{0,\mathrm{max}}(t=0)$ and $\rho_{\mathrm{cut,hi}}=\displaystyle 10~{}\rho_{\mathrm{cut,low}}$.
@@
Figure 10: A $xy$-plane slice of the star from the (S) simulation with the hybrid DG+FV/FD method at two times. Left: $t=0M_{\odot}$; right: $t=1000M_{\odot}$. We see that only the star surface continues to evolve with FV/FD at $t=1000M_{\odot}$, showing proper TEI functioning. The TEI adapts to the slight expansion of the star surface.

### v2

@@ IV.2 Special relativistic blast wave tests
As a next step, we propagated a so-called special relativistic blast wave in one dimension. The initial setup has discontinuities in the matter fields of $\rho_{0}$ and $P$, as in a Riemann problem. We study the case from [122] with $(\rho_{0},P,v_{x})$ given by
@@
We set up the mesh with $16$ adjacent Cartesian elements along the $x$-axis. Initially, we set the element at the left boundary and the middle element containing the initial discontinuity to FV/FD containing $24$ equally spaced grid points, and the rest of the elements to DG, with $12$ grid points. During the evolution, we use our TEI from Sec. III.7 to decide if an element should use the DG or FV/FD method. For this case, we switch off the criterion from Eq. (50), as the purpose of it is to handle low-density regions present in neutron star simulations, which is not relevant in this simulation. We also switch off the RDMP criterion, as it activates in regions where the analytic solution has zero slope, due to small amplitude variations in the numerical solution. These small amplitude variations can also trigger the Persson indicator if not relaxed enough. So, for the condition from Eq. (52), we choose $\alpha_{N}^{\mathrm{evo}}=4$ for the DG elements and $\alpha_{N}^{\mathrm{evo}}=5$ for the FV/FD elements. These values are lower than what we use for our star simulations in Secs. IV.3.2, IV.5.3 and IV.6, where we choose more conservative ones that result in more elements using FV/FD.
In Fig. (4) we plot the primitive variables. We show the entire mesh (top plot) as well as the contact discontinuity and the shock front (bottom plot), after evolving until time $t=0.4$. We see that the elements near the nonsmooth regions evolve with FV/FD, while most other elements evolve with DG, as can be seen from the top plot. In both plots, we see that the fields in our simulation agree very closely with the analytic solutions that are plotted as solid black lines. This demonstrates that our hybrid DG+FV/FD method can successfully handle evolutions of shock fronts in matter fields, which is a crucial initial test, as such hydrodynamic shocks are known to develop during late inspiral and merger phases of a binary neutron star merger.
@@ IV.3.1 Pure DG (S) simulations
For the DG case, we set up the mesh using a cubical patch at the center, surrounded by six cubed sphere patches on all sides, which in turn have another layer of six cubed sphere patches forming a spherical shell covering them, resulting in total 13 patches. We then h-refined the central cubical patch successively four, five and six times, while refining all the other patches one time less in all the cases, resulting in $10240$, $81920$, and $655360$ total elements after h-refining. In each of these elements, we use $5\times 5\times 5$ grid points. The star is contained in the central cube with side length $32$, which is centered on the star, while the outermost boundary of the entire mesh is at 320. The resultant initial number of elements across the star for the $l=4,5$, and $6$ are stated in Table 2. We set the time steps for the different resolutions to be $0.1$, $0.05$ and $0.025$ respectively. For the artificial atmosphere, we set $f_{\mathrm{atmo,level}}=10^{-12}$, $f_{\mathrm{atmo,cut}}=1.01\times 10^{-12}$, $f_{\mathrm{atmo,max}}=10$, $\epsilon_{\mathrm{atmo,max}}=0.1$, and $(Wv)_{\mathrm{atmo,max}}=0.01$ as per the description in Sec. III.5.1. We also apply our positivity limiters with elementwise limiting of $D$, $\tau$, and $S_{i}$, but not the pointwise fixes described in Sec. III.5.2.
@@
Figure 6: This figure shows the Fourier spectrum of the central point density $\rho_{0,\mathrm{C}}$ for (S) with both pure DG and DG+FV/FD and (P) with pure DG. The pure DG (S) and (P) data are taken from the simulations with $l=6$ levels of maximum h-refinement, while (S) DG+FV/FD has a maximum of $l=5$. The corresponding dimension of the finest elements resolving the star for the pure DG (S) and (P) cases is $738.313$ m, while that in the DG+FV/FD (S) case is $2399.516$ m. The average grid spacing in this finest element for the pure DG (S) and (P) simulations is $\Delta x_{\mathrm{avg,DG}}=184.578$ m. The DG elements in the interior of the star in the DG+FV/FD (S) case have $\Delta x_{\mathrm{avg,DG}}=342.788$ m, while FV/FD elements on the star surface have $\Delta x_{\mathrm{FV/FD}}=171.394$ m. The dashed vertical lines correspond to the expected frequencies for the various oscillations modes of the star, as stated in [123]. We see that the frequency for the principal and secondary modes from the oscillations matches perfectly with the theoretical value [123] for (S) with both pure DG and DG+FV/FD, while even the third peak matches for (S) with DG+FV/FD. The principal mode is also matching for the pure DG (P) case.
@@
This limiter treatment is essentially the same as that described in [76]. We restrict the lower limit of $D$ to $10^{-12}\rho_{0,\mathrm{max}}(t=0)$ and $\tau$ to 0, while enforcing $S<D+\tau$, as described when limiting. For the matter fields, as per Eq. (44), the filter uses $s=32$, while we use $s=40$ for the GH system. For both cases we have $\alpha_{f}=36$. For these simulations, we do not use the conserved to primitive variable conversion prescription mentioned in Sec. III.6. We retain the method for conversion described in Sec. 4.3.1 of [76]. This is because we performed simulation for the same physical system successfully under Cowling approximation with this method that has been presented in Sec. 4.3.4 in [76]. For the GH system we set the constraint damping parameters to be $\gamma_{0}=0.1$, $\gamma_{1}=-1$ and $\gamma_{2}=1$. The gauge is set such that the gauge source function is initialized according to the initial metric and then is kept constant throughout the simulation, i.e, we set
@@
Figure 7: This plot shows three quantities for (S) with pure DG for resolutions $l=4,5,$ and 66: total rest mass $\int DdV$ in the entire mesh normalized to the initial value (left), L2 norm of error in $D$ (center), and L2 norm of the Hamiltonian constraint (right). The error quantities (center and right) converge with order $2.6$. Mass conservation improves even faster with resolution. Here $dV=\sqrt{\gamma}d^{3}x$. The legend applies to all plots.
@@
The run with the five times h-refined central cube was evolved stably until a time of $t=5000M_{\odot}$, even though the plots for this case only show data until $t=2000M_{\odot}$ time due to high computational costs. As can be seen from Fig. 5, the central density oscillates with time, with higher erratic oscillations towards the beginning. These errors are larger in the beginning in this pure DG simulation due to strong Gibbs oscillations arising from the surface of the star. This is because in the initial setup, the nonsmooth feature on the surface of the star is at its sharpest, which produces large Gibbs oscillation in the DG scheme, as the surface lies in the middle of elements. These oscillations spread throughout the star and produce strong numerical disturbances. However, as the simulation progresses, our stabilizing mechanisms such as the positivity limiters, the filters and the atmosphere damp the spurious Gibbs oscillations, which leads to smoother star oscillations. Hence, at later times, we see a systematic oscillation for the two higher resolutions, which is in agreement with the spectrum expected from such a setup, as can be seen from Fig. 6, that shows the Fourier spectrum of the central density $\rho_{0,\mathrm{C}}(t)$ for the case with six central refinement levels for (S) with pure DG. For this simulation, the dimension of the finest elements covering the star is 0.5(738.31250.5~(738.3125 m), and the spacings between the Gauß-Lobatto grid points in these elements are: $\Delta x_{\mathrm{min,DG}}=0.0732~(108.123$ m), $\Delta x_{\mathrm{max,DG}}=0.177~(261.033$ m), and $\Delta x_{\mathrm{avg,DG}}=0.125~(184.578$ m), with the average being the grid spacing in an element of the same size with same number of equidistant grid points. In this figure we see that the principal frequency, as well as the second peak in the spectrum match the predicted value from literature [123]. The higher modes are buried in numerical noise due to the lack of resolution, though the third peak can be discerned to be approximately aligning with the correct value. Note that the numerical noise in the oscillations decreases with increasing resolution. We see the highest resolution is demonstrating consistent and persistent oscillations at the proper frequency, while for the middle resolution, they get damped slowly by the filters we apply. The lowest resolution of four central refinement levels is swamped by Gibbs oscillations, and is depicted only to show improvement with resolution. We next consider mass conservation by looking at the volume integral of the rest mass density $D$ shown in the left plot of Fig. 7. It is very clear that mass conservation improves drastically with increasing resolution. The small amount of discrepancy in the highest resolution is attributed to the fact that we use positivity limiters that conserve the mass only when the element average is above the limits imposed. Furthermore, pointwise corrections to $D$ arise due to the atmosphere treatment. In this case, the $D$ value is pushed up whenever it wants to fall below the floor value set by us. Hence, in these situations the integral can increase.
@@
We have also performed strong scaling tests with this (S) setup with pure DG, to demonstrate Nmesh’s capability to scale efficiently on a large number of computational cores or processors. This is shown in Fig. 8. The top plot shows speedup vs. number of cores used for the run, while the lower plot shows the parallel efficiency. Speedup when using $N$ number of compute cores is defined as $S(N)=t_{r}(1)/t_{r}(N)$, where $t_{r}(N)$ is the run-time measured for $N$ cores, with each core running one MPI process. The parallel efficiency is then defined as $E(N)=S(N)/N$, which is ideally $1$. Since a single MPI process cannot obtain enough memory for the simulations, $t_{r}(1)$ is estimated from $t_{r}(1)=N_{\mathrm{min}}t_{r}(N_{\mathrm{min}})$, where $N_{\mathrm{min}}$ is for the run with the lowest number of MPI processes performed in each case.
We performed these runs on both the Bridges-2 and Expanse supercomputers. In our case, we have $N_{\mathrm{min}}=3$ for the Expanse supercomputer and $512$ for the Bridges-2 supercomputer, which is used for the data in Fig. 8. We go up to $4096$ cores on Expanse and $6400$ cores on Bridges-2, which were the maximum core count available to us in the respective computers. The simulation we tested this with had total $262144$ elements in the mesh. We plot both the real simulation run-time data and the ideal ones for the number of cores we have in each of the cases, and show the speedups for both. The ideal speedups are calculated from the $t_{r}(N_{\mathrm{min}})$, as is the usual custom in such scaling estimates. We can see that we obtain good strong scaling, as the real speedups not only rise almost in a straight line, but track the ideal lines very closely. The fact that they overshoot somewhat is because we have only an estimate of $t_{r}(1)$ from $t_{r}(N_{\mathrm{min}})$. This is the same reason why the efficiency curve in the lower plot rises above $1$. This confirms that Nmesh scales efficiently with an increasing number of cores or MPI processes.
Figure 8: This plot shows speedup and parallel efficiency data for strong scaling tests on Bridges-2 and Expanse for the pure DG (S) simulation with $262144$ elements. The dashed lines in the top plot represent ideal speedup curves for the range of cores used on each supercomputer, while the solid lines represent the actual measured data for Nmesh. As we can see, the ideal and real data are in good agreement for speedup, demonstrating strong scaling.
IV.3.2 DG+FV/FD hybrid (S) simulation
@@
We also evolve (S) with our hybrid DG+FV/FD scheme. In the elements that use FV/FD, we use the WENOZ2 reconstruction scheme as described in Sec. III.2. Before using combined the DG+FV/FD method, we have performed tests with pure FV/FD with the WENOZ2 to validate our new method. However, we chose case (U) for these instead of (S), as it is significantly more challenging, and hence offers much more robust and general conclusions. These results are presented in Secs. IV.5.1 and IV.5.2. For the WENOZ2 method in this case, we turn on extrapolation near the boundary as per the $w=4/3$ setup for Eq. (III.2) described in Sec. III.2. Hence, we also switch on the cutoff levels for using $w=1$ as when $\rho_{0}<10^{-9}\rho_{0,max}(t=0)$ or $\epsilon<10^{-9}\epsilon_{max}(t=0)$, as described there. When the surface points of neighboring elements do not coincide, we use parabolic interpolation to exchange data between neighboring elements at the boundary. In the FD scheme used for the GH system we use fourth-order dissipation in the interior of the elements [see Eq. (33)] with $\sigma=0.1$. Since we want the high-order WENOZ used in the interior of elements to dominate the result, we increase the number of points in each direction in each FV/FD element to $n=16$, while in the DG elements, we use half as many points ($n=8$), as DG can provide high accuracy with lower number of points in the smooth regions. Each element then has $n\times n\times n$ total points. For this simulation, we set up the mesh similar to what we use for case (B), which is described later in Sec. IV.6, as they are the same star, with and without an imparted boost. The detail of the initial mesh can be seen from Fig. 9, which shows plots on the $xy$-plane. In the left top plot of the figure, we see the zoomed in view of the star. Initially, we set FV/FD only at the troubled region at the edge of the star. In order to determine which elements should be set to FV/FD at the initial time, we look at the coefficients of $\rho_{0}$ in an expansion in Legendre polynomials, as described in Sec. 2.5 of [76]. As we go from lower to higher order coefficients, we find a much slower dropoff for elements that contain the star surface, than for elements in the star interior. We use this dropoff to initially set only the elements containing the star surface to use FV/FD. These are recognizable from the denser grid lines as FV/FD uses twice as many points as DG. The initial number of DG and FV/FD elements across the star is given in Table 2. Next, in the top right plot we see the central cubical region, which is the best resolved region of the mesh. Here the size of this cube is $62$. However, as we can see, the resolution is not uniform. This is because we do not apply uniform h-refinement to the entire mesh. We h-refine a spherical region of radius $15$ centered on the star to $l=5$ times, which is the highest refined region in the mesh. The h-refinement is then dropped by one level as we move away from the star, with the different h-refinement spheres having radii $15,30,60,$ and $120$. These successive regions can be seen in the plots on the top right, bottom left, and bottom right. In the left bottom picture, we see the mesh up to the first cubed sphere patches surrounding the central cube, up to a distance of $104$. The last bottom right picture shows the entire mesh, which has a radius of $208$. In total, we have 13 patches similar to the setup in Sec. IV.3.1. In the most refined central region (containing the star) in this arrangement, the elements have a dimension of 1.625(2399.5161.625~(2399.516 m). Of these, the DG elements in the interior of the star have the following grid spacing: $\Delta x_{\mathrm{min,DG}}=0.0805~(118.813$ m), $\Delta x_{\mathrm{max,DG}}=0.362~(533.943$ m), and $\Delta x_{\mathrm{avg,DG}}=0.232~(342.788$ m). On the other hand, the FV/FD elements on the star surface have uniform grid spacing $\Delta x_{\mathrm{FV/FD}}=0.116~(171.394$ m). This is because we have twice the number of grid points in the FV/FD elements compared to the DG elements. The time step is set as per Sec. III.3.
@@
For the atmosphere we use the values $f_{\mathrm{atmo,level}}=10^{-11}$, $f_{\mathrm{atmo,cut}}=10^{-9}$, and $f_{\mathrm{atmo,max}}=1$, as this is the setup we also use for (U) and (B) simulations with DG+FV/FD, which are described later. The other atmosphere limits are still maintained at $\epsilon_{\mathrm{atmo,max}}=0.1$, and $(Wv)_{\mathrm{atmo,max}}=0.01$. Here, we employ the entire limiting process described in Sec. III.5.2 for $D$, $\tau$, and $S_{i}$, including both elementwise and pointwise limiting. For this purpose, we set $f_{S}^{\mathrm{lim}}=0.99$. In this simulation, we use the conversion procedure from conserved to primitive variables stated in Sec. III.6, setting $\epsilon_{\mathrm{floor}}=0.0$. The constraint damping parameters for the GH system are set to be the position dependent functions of the form from Eq. (5) along with Eqs. (59)-(61). The gauge is chosen to be the harmonic gauge, where the gauge source function is set as $H_{\alpha}=0$. In the DG elements, we filter the conserved matter variables $D,\tau$, and $S_{i}$ with $s=40$, while using $s=64$ for the GH variables of the evolved metric and its derivatives. We maintain $\alpha_{f}=36$ for both. No filters are used in the FV/FD elements, as dissipation plays the counterpart role. When incorporating the TEI described in Sec. III.7, we set the value of $\alpha_{N}^{\mathrm{evo}}$ from Eq. (52) in the elements using DG to be $8$ and $9$ in the elements using FV/FD. For this criterion, we only consider the coefficients that are greater than $f_{\mathrm{unfiltered}}=0.9$ times their unfiltered value. We set the cutoff from Eq. (50) as $\rho_{\mathrm{cut,low}}=10^{-4}\rho_{0,\mathrm{max}}(t=0)$ and $\rho_{\mathrm{cut,hi}}=\displaystyle 10~\rho_{\mathrm{cut,low}}$.
@@
Figure 10: An xyxy-plane slice of the star from the (S) simulation with the hybrid DG+FV/FD method at two times. Left: $t=0M_{\odot}$; right: $t=1000M_{\odot}$. We see that only the star surface continues to evolve with FV/FD at $t=1000M_{\odot}$, showing proper TEI functioning. The TEI adapts to the slight expansion of the star surface.